OpenStudy (anonymous):

If x =3 and y = -2 , then y4 - 4x = ? A. -28 B. -20 C. -4 D. 4

OpenStudy (anonymous):

Substitute the values into the equation and evaluate.

OpenStudy (anonymous):

is it A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

D.4 is the ans

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

\[\Large - 2 \cdot 4 - 4 \cdot 3 = - 8 - 12 = - 20\]

OpenStudy (anonymous):

isn't it y^4

OpenStudy (anonymous):

Possibly... correct notation would be helpful :-P If it's y^4 then...\[\Large - {2^4} - 4 \cdot 3 = 16 - 12 = 4\]